A building contractor gives a ​$14,000 promissory note to a plumber who has loaned him ​$14,000. The note is due in 9 months with interest at 7​%. Three months after the note is​ signed, the plumber sells it to a bank. If the bank gets a 9​% return on its​ investment, how much will the plumber​ receive? Will it be enough to pay a bill for ​$14,066​? How much will the plumber​ receive?
Answer #1

Note: We are assuming that the interest rates given in question are per annum interest rate

Promissory note value after 9 months =14000+ (14000*7%*9/12 )=14000+735 = $14735

The plumber sells the note to bank after 3 months so remaining period for note maturity = 9-3 = 6 months

FV = PV(1+r)^n

We know the FV of note as calculated above i.e. 14735

Rate of interest charged by bank for 6 months = 9/2 = 4.5%

Therefore, 14735 = PV(1+0.045)

PV = 14735/1.045 = 14100.478469

Therefore, if the plumber sells the note to bank he will receive $14100.478469

Since plumber will receive $14100.478469 it will be enough to pay the bill for $14066